# Numerical simulation and analysis of mixing enhancement due to chaotic advection using an adaptive approach for approximating the dilution index

### v1

@@ 3.5 A computationally reliable approach for the approximation of the dilution index
Figure 8: (a)-(b) Reactor ratio $M$ and (c)-(d) reactor ratio growth rate over time $t$ for the PSS system with different values of $\Lambda$. The growth rate was approximated using the moving average over the first order finite differences with a window of ±$\pm 0.05$ in the exponent. The dotted lines correspond to the case without advection where the initial particle position is either at one point (gray) or in a ring (black).
Figure 9: (a)-(b) Reactor ratio $M$ plotted with a logarithmically typed $y$-axis highlighting the convergence $M\rightarrow 1$ and (c)-(d) convergence rate of the reactor ratio for different values of $\Lambda$ in the PSS system computed using (20). The convergence rate was approximated with the moving average over the first order finite differences with a window of ±$\pm 0.05$ in the exponent. The dotted lines correspond to the case without advection where the initial particle position is either at one point (gray) or in a ring (black). Note that for $\sigma=0.01\beta$ the difference between the results of point and the ring initialization is below the resolution of the plot.
@@ 4.1 Pulsed source-sink
We expect the reactor ratio to increase monotonically over time while approaching $M_{\text{max}}=1$. We also expect faster dilution in the chaotic advection setups compared to the baseline experiments. In Fig. 7, it is possible to see the effect of chaotic advection on the evolution of the probability density of the particle distribution for $\Lambda^{2}=0.2$. For this and all other considered flow parametrization, we show the approximated reactor ratio $M$ over time $t$ in Fig. 8(a) and (b). Figures 8(c) and (d) show the reactor ratio growth rate, which is always positive, demonstrating that our method preserves the monotonical increase. The baseline cases without advection display constant growth rates for small times. For the point initialization, the plume is distributed according to the Gaussian distribution with variance $\sigma^{2}t$ for small $t$. The reactor ratio is therefore given by $M(t)=\frac{\pi}{8}\exp(1)\sigma^{2}t$, which can only increase linearly in time. The ring initialization grows at a constant rate of $0.5$ for small times. The reason is that diffusion can only act in the direction perpendicular to the ring due to the ring structure. As shown in Kitanidis (1994), the dilution index of the Gaussian distribution in one dimension with variance $\sigma^{2}t$ is given by $\sqrt{2\pi t}\,\sigma^{2}\exp\left(\frac{1}{2}\right)$ which has a growth rate of $0.5$ in time. For later times, the particles fill the ring and appear more similar to a two-dimensional Gaussian distribution. Therefore, the growth rate can reach up to one, and for later times and $\sigma=0.1\beta$, the reactor ratios of both baseline experiments coincide approximately. Despite the lower growth rate, we note that the reactor ratio for the baseline with ring initialization is higher than for the point initialization. This behavior is expected due to the higher entropy of the particle location. In the chaotic advection systems, the growth rates for small times are also close to $0.5$. At the beginning, the particles form a circle around the source, and thus, the behavior is similar to the baseline with the ring initialization. As the particles move away from the source, the growth rate increases slowly until $t\approx 0.6$ when the growth rate experiences a remarkable rise due to the first cluster of particles reaching the sink. As we chose $Q$ to be the same for all configurations, the particles arrive at the sink almost simultaneously, explaining the simultaneous rise of the growth rates for all experiments, i.e., for different choices of $\Lambda^{2}$ and $\sigma$. Due to the boundedness of the reactor ratio, the growth rate cannot remain at a high level for a long time. The rate quickly falls to zero, especially for $\sigma=0.1\beta$ (Fig. 8(d)). As the particles fill the domain in Fig. 7 bottom $M$ approaches $M_{\text{max}}$. However, this already shows the increased dilution caused by chaotic advection. Considering now Fig. 9(a) and (b), which also show the graph $M$ over $t$, but the axis for the reactor ratios have a logarithmic scaling to highlight how $M$ approaches $1$. For larger times and $\sigma=0.1\beta$, the reactor ratio reaches a maximum degree of dilution, which is strictly smaller than $1$. This effect is due to low particle density regions between domain boundary and source/sink also observed in the Poincaré sections in Fig. 3, which is an artifact of the periodic-like boundary condition. Looking at the convergence rates of the reactor ratio, approaching $M_{\text{max}}$ in Fig. 9(c) and (d) and represent the visible derivative of the curves in Fig. 9(a) and (b). The convergence rates at time $t$ can be computed numerically via
@@ 4.2 RPM flow
Figure 10: (a)-(b) Evolution of $M$ over $t$ for the RPM flow with a logarithmically typed $y$-axis and (c)-(d) the convergence rate for $M$ approaching $1$ for different flow parametrization computed using (20). The convergence rate was computed using the moving average over the first order finite difference by using a window of ±$\pm 0.2$ in the exponent. The dotted lines correspond to the case without advection where the initial particle position is either at one point (gray) or in a ring (black). Note that for $\sigma=0.01$ the difference between the results of point and the ring initialization is below the resolution of the plot.
@@
During the RPM system experiment, $1$ million particles are initially placed at the location of the source. This initialization represents that they are injected through the source at $t=0$. The parameters are chosen as motivated above with the rotation angle $\Theta\in\{0,\frac{\pi}{6},\frac{\pi}{3},\frac{\pi}{2},\frac{2\pi}{3},\frac{5\pi}{6}\}$, the time between rotations $\tau\in\{0.2,0.5\}$, the diffusion coefficient $\sigma\in\{0.01,0.1\}$, $t_{\text{max}}=20$ and $\Delta t=0.01$. While $\Theta>0$ displays chaotic advection, we do not find chaotic flow structures for $\Theta=0$. However, we observe shear flow and focusing and defocusing flow, which also leads to deformation of the plume and, therefore, to increased mixing Rolle and Le Borgne (2019); Paster et al. (2015). Like for the pulsed source-sink system, we consider two baseline setups without advection, but with the same reflective boundary conditions as for the systems with advection. We also use the point initialization, where all particles are initially placed at the origin, and the ring initialization, where we take the locations of the particles of the advection experiments at time $t=0.01$ as initialization. When approximating the dilution index, we estimate the grid size $h$ for each considered time step individually, by using the minimum of $\Delta_{\delta}[E](h)$ defined in (19), ensuring that $h$ is monotonically increasing in time and $h\leq 0.1$.
@@
Figure 12: Evolution of the probability density of the particle distribution over time using $(\Theta,\tau)=\left(\frac{\pi}{2},0.5\right)$ and $\sigma\in\{0.01,0.1\}$ for the RPM flow.
@@
In Fig. 10(a) and (b) we show the evolution of the reactor ratio $M$ over $t$ for $\Theta\in\left\{0,\frac{\pi}{6},\frac{\pi}{2}\right\}$ and $\tau\in\{0.2,0.5\}$. The $y$-axis has a logarithmic typed rescaling to highlight the convergence $M\rightarrow 1$. For smaller times, the reactor ratio increases monotonically, as expected. For $\sigma=0.1$ and large times, some configurations reach a status of full dilution. Due to the finite amount of particles, the approximated density distribution always differs slightly from the uniform distribution; therefore, the reactor ratio must be smaller than $1$, and we observe oscillations resulting in occasional decreases in $M$. Comparing the dilution enhancement of different flow parameter configurations, we notice significant differences. Investigating this further, Fig. 11(a) and Fig. 11(b) show the probability density approximations and $M$ over $\Theta$, respectively, at time $t=10$, where reactor ratios differ remarkably between the configurations. Each figure column represents a rotation angle $\Theta$, and Fig. 11(a) includes the cases without advection on the right. For $\sigma=0.01$ and $\Theta>0$, the KAM islands found in the Poincaré maps of Fig. 4 are mainly particle-free. This is due to the periodic nature of the points inside the islands. As the particles are not initially placed on trajectories leading through these islands, they can only enter due to the diffusion process. This makes the dilution inside these regions far slower than in the chaotic area of the domain. For $\Theta=0$, no KAM islands arise in the flow, but due to the non-chaotic nature of the flow, particles fill the domain much slower. We, therefore, achieve faster dilution for chaotic flow fields with only a few small KAM islands as for $(\Theta,\tau)=\left(\frac{\pi}{6},0.5\right)$ or $\left(\frac{\pi}{2},0.2\right)$. In general, this implies that the non-chaotic configuration $\Theta=0$ is not necessarily worse than $\Theta>0$ in terms of mixing, in particular when compared with configurations leading to big KAM islands. This effect does not occur in the pulsed source-sink system where all trajectories, including the periodic ones, pass through the sink and the source. By initially placing the particles on the source, some of them end up on the periodic trajectories and fill the KAM islands.
@@
To quantify the effect of KAM islands, we introduce a quasi reactor ratio. To do so, we simulate the motion of particles with $\sigma=0$ and approximate the size of the chaotic region. We use a grid with size $h=0.01$ and count all cells with at least one particle. The quasi reactor ratio $\bar{M}(\sigma=0)$ is given as the fraction of the chaotic area to the full domain (black lines in Fig. 11(b)). We see that the reactor ratios approximately follow the same trend as the quasi reactor ratio. The parameter configuration $\left(\Theta,\tau\right)=\left(\frac{\pi}{6},0.5\right)$ leads to the highest quasi reactor ratio and corresponds to the configuration with the fastest dilution.
@@
Figure 13: (a)-(b) Reactor ratio $M$ and (c)-(d) reactor ratio growth rate over time $t$ for different flow parametrization for the RPM flow. The growth rate was approximated using the moving average over the first order finite differences with a window of ±$\pm 0.2$ in the exponent. The dotted lines correspond to the baseline cases without advection where the initial particle position is either at one point (gray) or in a ring (black).
@@
Summarizing the above, it appears that the individual growth of the reactor ratio is very dependent on the KAM island structure. Therefore, the growth rates and convergence rates of the reactor ratio vary over the parameter space. Looking at the convergence rates in Fig. 10(c) (computed using (20)) for $\sigma=0.01$, we find that they reach a maximum before the rate decreases again. To understand this behavior we look at the temporal evolution of the probability density distributions of the particles evolved using $\Theta=\frac{\pi}{2}$ and $\tau=0.5$ in the top row of Fig. 12. At the start of the time interval of high convergence rates $t\approx 2$, the particles are concentrated at a few stripes in the domain. From here, they quickly spread until the particles are well diluted inside the chaotic area. To also reach high dilution inside the KAM island, particles need to diffuse into the islands. Since in advective dominated flows diffusive transport is more time-consuming than advective transport the convergence rates of $M\rightarrow 1$ are lower after reaching high dilution in the chaotic area. For a higher diffusion coefficient (e.g., $\sigma=0.1$), entering the KAM islands is not a time-consuming process (Fig. 12, bottom). For later times, we observe convergence rates close to zero, which is a strong indicator of a setup to be fully diluted. These results demonstrate that diffusion plays an important role in systems with chaotic advection, as it is the limiting process for filling of the KAM islands.
@@
Finally, we consider the growth rates of $M$ in Fig. 13 and compare the results of the advective setups to the growth rates of the setups without advection. As the first rotation happens at time $t=\tau$, the graphs coincide for $t<\tau$, and the growth rate for small times is close to $0.5$ and increases in time. After some rotations of source and sink, the growth rates may eventually start rising above $1$, especially for $\sigma=0.01$, where most of the setups reach a maximal growth rate of around $2$ (Fig. 13(c)). At higher diffusivity, such as $\sigma=0.1$ (Fig. 13(d)), the reactor ratio gets close to the maximum much earlier, resulting in lower maximal growth rates. Unlike for the pulsed source-sink experiment, the dilution indices of the baseline experiment for $\sigma=0.1$ do not coincide for later times. This is the effect of the reflective boundary condition, which decreases the effect of diffusion close to the domain boundary.
@@ 5 Conclusion
In this work, we present two contributions. First, we develop an accurate method to adaptively select a grid size for the approxmation of an entropy-based measure, the dilution index, preserving its key feature to monotonically increase in time. Our approach allows the dilution index to effectively quantify the enhanced mixing between a solute and a solvent in chaotic flow fields. In the second part, we apply our method to analyze two chaotic injection-extraction systems, revealing that not all chaotic setups lead to very high mixing enhancement.
@@
Evaluating the dilution index to assess the results of particle tracing simulations reveals an undesirable result. The approximated dilution index can yield any value in $(0,E_{\text{max}}]$, depending on the selected grid size, but independent of the underlying particle distribution. Given a sufficient amount of particles, we provide a fixed criterion on the grid size motivated by the theory of representative elementary volumes. Our method is the first approach to sophisticatedly selecting the grid size of the dilution index to evaluate particle tracking experiments, guaranteeing accurate and unbiased results that show monotonic growth over time. Increasing the number of particles enables the choice of smaller grid sizes and improves the accuracy of the approximation.
@@
We apply our method to evaluate the temporal increase in mixing of a solute within its solvent due to chaotic advection. Our analysis shows that only some configurations of chaotic advection can enhance mixing compared to non-chaotic advection, as KAM islands limit the chaotic region of the chaotic flow. The results demonstrate the fundamental role of diffusion in systems with chaotic advection in enhancing mixing. Diffusion significantly contributes to filling the KAM island and, hence, reaching complete mixing in the systems. Comparing to a baseline without advection, configurations with advection consistently lead to faster mixing. Our analysis does not consider the effect of the initial solute particle location. In a low diffusivity setup where the solute is initially placed inside a KAM island, the mixing in the chaotic setups can be worse than that of non-chaotic advection.
@@
Considering chaotic injection-extraction systems for groundwater remediation purposes, our research emphasizes the importance of an analysis of the mixing potential of a system prior to its implementation into the field. Due to high uncertainties in the plume location, a configuration with small KAM islands is preferred to achieve the best remediation results. In order to generate more accurate predictions on the expected mixing enhancement in the field, we envision future research considering heterogeneities in the domain as they would appear due to the geological characteristics of the soil. As it is shown that these soil heterogeneities lead to increased mixing, we expect interactions with the chaotic flow field, which will modify the KAM island structure. We can further reduce KAM islands by randomizing the flow fields. If the flow parameters of the chaotic setups are varied randomly in time, the KAM island structure may disappear, increasing the mixing potential.

### v2

@@ 3.5 A computationally reliable approach for the approximation of the dilution index
Figure 8: (a)-(b) Reactor ratio $M$ and (c)-(d) reactor ratio growth rate over time $t$ for the PSS system with different values of $\Lambda$. The growth rate was approximated using the moving average over the first order finite differences with a window of $\pm 0.05$ in the exponent. The dotted lines correspond to the case without advection where the initial particle position is either at one point (gray) or in a ring (black).
Figure 9: (a)-(b) Reactor ratio $M$ plotted with a logarithmically typed $y$-axis highlighting the convergence $M\rightarrow 1$ and (c)-(d) convergence rate of the reactor ratio for different values of $\Lambda$ in the PSS system computed using (20). The convergence rate was approximated with the moving average over the first order finite differences with a window of $\pm 0.05$ in the exponent. The dotted lines correspond to the case without advection where the initial particle position is either at one point (gray) or in a ring (black). Note that for $\sigma=0.01\beta$ the difference between the results of point and the ring initialization is below the resolution of the plot.
@@ 4.1 Pulsed source-sink
We expect the reactor ratio to increase monotonically over time while approaching $M_{\text{max}}=1$. We also expect faster dilution in the chaotic advection setups compared to the baseline experiments. In Fig. 7, it is possible to see the effect of chaotic advection on the evolution of the probability density of the particle distribution for $\Lambda^{2}=0.2$. For this and all other considered flow parametrization, we show the approximated reactor ratio $M$ over time $t$ in Fig. 8(a) and (b). Fig. 8(c) and (d) show the reactor ratio growth rate, which is always positive, demonstrating that our method preserves the monotonical increase. The baseline cases without advection display constant growth rates for small times. For the point initialization, the plume is distributed according to the Gaussian distribution with variance $\sigma^{2}t$ for small $t$. The reactor ratio is therefore given by $M(t)=\pi/8\exp(1)\sigma^{2}t$, which can only increase linearly in time. The ring initialization grows at a constant rate of $0.5$ for small times. The reason is that diffusion can only act in the direction perpendicular to the ring due to the ring structure. As shown in [20], the dilution index of the Gaussian distribution in one dimension with variance $\sigma^{2}t$ is given by $\sqrt{2\pi t}\,\sigma^{2}\exp\left(1/2\right)$ which has a growth rate of $0.5$ in time. For later times, the particles fill the ring and appear more similar to a two-dimensional Gaussian distribution. Therefore, the growth rate can reach up to one, and for later times and $\sigma=0.1\beta$, the reactor ratios of both baseline experiments coincide approximately. Despite the lower growth rate, we note that the reactor ratio for the baseline with ring initialization is higher than for the point initialization. This behavior is expected due to the higher entropy of the particle location. In the chaotic advection systems, the growth rates for small times are also close to $0.5$. At the beginning, the particles form a circle around the source, and thus, the behavior is similar to the baseline with the ring initialization. As the particles move away from the source, the growth rate increases slowly until $t\approx 0.6$ when the growth rate experiences a remarkable rise due to the first cluster of particles reaching the sink. As we chose $Q$ to be the same for all configurations, the particles arrive at the sink within a similar time frame, explaining the simultaneous rise of the growth rates for all experiments, i.e., for different choices of $\Lambda^{2}$ and $\sigma$. Due to the boundedness of the reactor ratio, the growth rate cannot remain at a high level for a long time. The rate quickly falls to zero, especially for $\sigma=0.1\beta$ (Fig. 8(d)). As the particles fill the domain in Fig. 7 bottom $M$ approaches $M_{\text{max}}$. However, this already shows the increased dilution caused by chaotic advection. Considering now Fig. 9(a) and (b), which also show the graph $M$ over $t$, but the axis for the reactor ratios have a logarithmic scaling to highlight how $M$ approaches $1$. For larger times and $\sigma=0.1\beta$, the reactor ratio reaches a maximum degree of dilution, which is strictly smaller than $1$. This effect is due to low particle density regions between domain boundary and source/sink also observed in the Poincaré sections in Fig. 3, which is an artifact of the periodic-like boundary condition. Looking at the convergence rates of the reactor ratio, approaching $M_{\text{max}}$ in Fig. 9(c) and (d) and represent the visible derivative of the curves in Fig. 9(a) and (b). The convergence rates at time $t$ can be computed numerically via
@@ 4.2 RPM flow
Figure 10: (a)-(b) Evolution of $M$ over $t$ for the RPM flow with a logarithmically typed $y$-axis and (c)-(d) the convergence rate for $M$ approaching 11 for different flow parametrization computed using (20). The convergence rate was computed using the moving average over the first order finite difference by using a window of $\pm 0.2$ in the exponent. The dotted lines correspond to the case without advection where the initial particle position is either at one point (gray) or in a ring (black). Note that for $\sigma=0.01$ the difference between the results of point and the ring initialization is below the resolution of the plot.
@@
During the RPM system experiment, $1$ million particles are initially placed at the location of the source. This initialization represents that they are injected through the source at $t=0$. The parameters are chosen as motivated above with the rotation angle $\Theta\in\left\{0,\pi/6,\pi/3,\pi/2,2\pi/3,5\pi/6\right\}$, the time between rotations $\tau\in\{0.2,0.5\}$, the diffusion coefficient $\sigma\in\{0.01,0.1\}$, $t_{\text{max}}=20$ and $\Delta t=0.01$. While $\Theta>0$ displays chaotic advection, we do not find chaotic flow structures for $\Theta=0$. However, we observe shear flow and focusing and defocusing flow, which also leads to deformation of the plume and, therefore, to increased mixing [38, 34]. Like for the pulsed source-sink system, we consider two baseline setups without advection, but with the same reflective boundary conditions as for the systems with advection. We also use the point initialization, where all particles are initially placed at the origin, and the ring initialization, where we take the locations of the particles of the advection experiments at time $t=0.01$ as initialization. When approximating the dilution index, we estimate the grid size $h$ for each considered time step individually, by using the minimum of $\Delta_{\delta}[E](h)$ defined in (19), ensuring that $h$ is monotonically increasing in time and $h\leq 0.1$.
@@
Figure 12: Evolution of the probability density of the particle distribution over time using $(\Theta,\tau)=\left(\pi/2,0.5\right)$ and $\sigma\in\{0.01,0.1\}$ for the RPM flow.
@@
In Fig. S2(a) and (b) we show the evolution of the reactor ratio $M$ over $t$ for $\Theta\in\left\{0,\pi/6,\pi/2\right\}$ and $\tau\in\{0.2,0.5\}$. Fig. S2 in section S3 of the supplementary material includes the results for all considered parameter combination. The $y$-axis has a logarithmic typed rescaling to highlight the convergence $M\rightarrow 1$. For smaller times, the reactor ratio increases monotonically, as expected. For $\sigma=0.1$ and large times, some configurations reach a status of full dilution. Due to the finite amount of particles, the approximated density distribution always differs slightly from the uniform distribution; therefore, the reactor ratio must be smaller than $1$, and we observe oscillations resulting in occasional decreases in $M$. Comparing the dilution enhancement of different flow parameter configurations, we notice significant differences. Investigating this further, Fig. 11(a) and Fig. 11(b) show the probability density approximations and $M$ over $\Theta$, respectively, at time $t=10$, where reactor ratios differ remarkably between the configurations. Each figure column represents a rotation angle $\Theta$, and Fig. 11(a) includes the cases without advection on the right. For $\sigma=0.01$ and $\Theta>0$, the KAM islands found in the Poincaré maps of Fig. 4 are mainly particle-free. This is due to the periodic nature of the points inside the islands. As the particles are not initially placed on trajectories leading through these islands, they can only enter due to the diffusion process. This makes the dilution inside these regions far slower than in the chaotic area of the domain. For $\Theta=0$, no KAM islands arise in the flow, but due to the non-chaotic nature of the flow, particles fill the domain much slower. We, therefore, achieve faster dilution for chaotic flow fields with only a few small KAM islands as for $(\Theta,\tau)=\left(\pi/6,0.5\right)$ or $\left(\pi/2,0.2\right)$. In general, this implies that the non-chaotic configuration $\Theta=0$ is not necessarily worse than $\Theta>0$ in terms of mixing, in particular when compared with configurations leading to big KAM islands. This effect does not occur in the pulsed source-sink system where all trajectories, including the periodic ones, pass through the sink and the source. By initially placing the particles on the source, some of them end up on the periodic trajectories and fill the KAM islands.
@@
To quantify the effect of KAM islands, we introduce a quasi reactor ratio. To do so, we simulate the motion of particles with $\sigma=0$ and approximate the size of the chaotic region. We use a grid with size $h=0.01$ and count all cells with at least one particle. The quasi reactor ratio $\bar{M}(\sigma=0)$ is given as the fraction of the chaotic area to the full domain (black lines in Fig. 11(b)). We see that the reactor ratios approximately follow the same trend as the quasi reactor ratio. The parameter configuration $\left(\Theta,\tau\right)=\left(\pi/6,0.5\right)$ leads to the highest quasi reactor ratio and corresponds to the configuration with the fastest dilution.
@@
Figure 13: (a)-(b) Reactor ratio $M$ and (c)-(d) reactor ratio growth rate over time $t$ for different flow parametrization for the RPM flow. The growth rate was approximated using the moving average over the first order finite differences with a window of $\pm 0.2$ in the exponent. The dotted lines correspond to the baseline cases without advection where the initial particle position is either at one point (gray) or in a ring (black).
@@
Summarizing the above, it appears that the individual growth of the reactor ratio is very dependent on the KAM island structure. Therefore, the growth rates and convergence rates of the reactor ratio vary over the parameter space. Looking at the convergence rates in Fig. S2(c) (computed using (20)) for $\sigma=0.01$, we find that they reach a maximum before the rate decreases again. To understand this behavior we look at the temporal evolution of the probability density distributions of the particles evolved using $\Theta=\pi/2$ and $\tau=0.5$ in the top row of Fig. 12. At the start of the time interval of high convergence rates $t\approx 2$, the particles are concentrated at a few stripes in the domain. From here, they quickly spread until the particles are well diluted inside the chaotic area. To also reach high dilution inside the KAM island, particles need to diffuse into the islands. Since in advective dominated flows diffusive transport is more time-consuming than advective transport the convergence rates of $M\rightarrow 1$ are lower after reaching high dilution in the chaotic area. For a higher diffusion coefficient (e.g., $\sigma=0.1$), entering the KAM islands is not a time-consuming process (Fig. 12, bottom). For later times, we observe convergence rates close to zero, which is a strong indicator of a setup to be fully diluted. These results demonstrate that diffusion plays an important role in systems with chaotic advection, as it is the limiting process for filling of the KAM islands.
@@
Finally, we consider the growth rates of $M$ in Fig. S3 and compare the results of the advective setups to the growth rates of the setups without advection. For the results including all considered parameter configurations we refer to the supplementary material Fig S3. As the first rotation happens at time $t=\tau$, the graphs coincide for $t<\tau$, and the growth rate for small times is close to $0.5$ and increases in time. After some rotations of source and sink, the growth rates may eventually start rising above $1$, especially for $\sigma=0.01$, where most of the setups reach a maximal growth rate of around $2$ (Fig. S3(c)). At higher diffusivity, such as $\sigma=0.1$ (Fig. S3(d)), the reactor ratio gets close to the maximum much earlier, resulting in lower maximal growth rates. Unlike for the pulsed source-sink experiment, the dilution indices of the baseline experiment for $\sigma=0.1$ do not coincide for later times. This is the effect of the reflective boundary condition, which decreases the effect of diffusion close to the domain boundary.
@@ 5 Conclusion
In this work, we present two contributions. First, we develop an accurate method to adaptively select a grid size for the approximation of an entropy-based measure, the dilution index, preserving its key feature to monotonically increase in time. Our approach allows the dilution index to effectively quantify the enhanced mixing between a solute and a solvent in chaotic flow fields. In the second part, we apply our method to analyze two chaotic injection-extraction systems, revealing that not all chaotic setups lead to very high mixing enhancement.
@@
Evaluating the dilution index to assess the results of particle tracing simulations reveals an undesirable result. The approximated dilution index can yield any value in $(0,E_{\text{max}}]$, depending on the selected grid size, but independent of the underlying particle distribution. Given a sufficient amount of particles, we provide a fixed criterion on the grid size motivated by the theory of representative elementary volumes. Our method is the first approach to sophisticatedly select the grid size of the dilution index to evaluate particle tracking experiments, guaranteeing accurate and unbiased results that show monotonic growth over time. Increasing the number of particles enables the choice of smaller grid sizes and improves the accuracy of the approximation.
@@
We apply our method to evaluate the temporal increase in mixing of a solute within its solvent due to chaotic advection. To approximate the solution of the advection-diffusion equation, we use lagrangian particle tracking to preserve the fine plume structures due to chaotic advection and minimize numerical diffusion. Our analysis shows that only some configurations of chaotic advection can enhance mixing compared to non-chaotic advection, as KAM islands limit the chaotic region of the chaotic flow. The results demonstrate the fundamental role of diffusion in systems with chaotic advection in enhancing mixing. Diffusion significantly contributes to filling the KAM island and, hence, reaching complete mixing in the systems. Comparing to a baseline without advection, configurations with advection consistently lead to faster mixing. Our analysis does not consider the effect of the initial solute particle location. In a low diffusivity setup where the solute is initially placed inside a KAM island, the mixing in the chaotic setups can be worse than that of non-chaotic advection.
@@
Considering chaotic injection-extraction systems for groundwater remediation purposes, our research emphasizes the importance of an analysis of the mixing potential of a system prior to its implementation into the field. Due to high uncertainties in the plume location, a configuration with small KAM islands is preferred to achieve the best remediation results. In order to generate more accurate predictions on the expected mixing enhancement in the field, we envision future research considering heterogeneities in the domain as they would appear due to the geological characteristics of the soil. As it is shown that these soil heterogeneities lead to increased mixing, we expect interactions with the chaotic flow field, which will modify the KAM island structure. To account for the non-Fickian transport in heterogeneous porous media, it would be necessary to adapt our particle tracking method according to [46]. The topic of randomizing the flow field is additionally important, as it can further reduce KAM islands. If the flow parameters of the chaotic setups are varied randomly in time, the KAM island structure may disappear, increasing the mixing potential.
CRediT authorship contribution statement
Carla Feistner: Conceptualization, Formal analysis, Investigation, Methodology, Software, Visualization, Writing – original draft, Mónica Basilio Hazas: Writing – review & editing, Barbara Wohlmuth: Supervision, Writing – review & editing, Gabriele Chiogna: Conceptualization, Funding acquisition, Project administration, Supervision, Writing – review & editing
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.